OpenStudy (anonymous):

simplify the square root of 4/9

OpenStudy (anonymous):

\[\sqrt{\frac{ 4 }{ 9 }}?\] or \[\frac{ \sqrt{4} }{ 9 }?\]

OpenStudy (anonymous):

the first one

OpenStudy (anonymous):

Do you have a calculator?

OpenStudy (anonymous):

no i do not at this moment

OpenStudy (anonymous):

Oh okay well I plugged it in and it came up as .66666666667 or known as 2/3

OpenStudy (anonymous):

ok thanks a lot for your reply